Write the slope-intercept form of the equation that is perpendicular to the line 6x-4y=8 and passes through the point (1, 4).
Answer by mananth(16949)

6 x -4 y = 8
Find the slope of this line

-4y=-6x + 8
Divide by -4
y=2x-2
Compare this equation with y=mx+b
slope m = 1.50=>3/2

The slope of a line perpendicular to the above line will be the negative reciprocal-2/3
m1*m2=-1
The slope of the required line will be - 2/3

m= - 2/3 ,point ( 1 , 4 )
Find b by plugging the values of m & the point in
y=mx+b
4 = -2/3 + b
b= 14/3
m= - 2/3
The required equation is y = - 2/ 3 x+ 14/ 3
m.ananth@hotmail.ca